OpenStudy (anonymous):

determine the slope of a line passing through the points (2, 9) and (0.66, 9)

OpenStudy (accessdenied):

Slope is given by: \[ \large m = \frac{y_2 - y_1}{x_2 - x_1} \] *

OpenStudy (anonymous):

We have to use the gradient and find an equation formula \[M=\frac{y2-y1}{x2-}\] <<gradient/slope formula \[y-y1 = m (x-x1)\] <<find an equation formula \[M=\frac{9-9}{0.66-2}= \frac{0}{-1.44} = 0\] now use the find an equation formula and sub in M=0 and one point (2,9) x1,y1

OpenStudy (anonymous):

thank you